# MultiMat: Multimodal Program Synthesis for Procedural Materials using Large Multimodal Models

### v1

@@ Abstract
Material node graphs are programs that generate the 2D channels of procedural materials, including geometry such as roughness and displacement maps, and reflectance such as albedo and conductivity maps. They are essential in computer graphics for representing the appearance of virtual 3D objects parametrically and at arbitrary resolution. In particular, their directed acyclic graph structures and intermediate states provide an intuitive understanding and workflow for interactive appearance modeling. Creating such graphs is a challenging task and typically requires professional training. While recent neural program synthesis approaches attempt to simplify this process, they solely represent graphs as textual programs, failing to capture the inherently visual-spatial nature of node graphs that makes them accessible to humans. To address this gap, we present MultiMat, a multimodal program synthesis framework that leverages large multimodal models to process both visual and textual graph representations for improved generation of procedural material graphs. We train our models on a new dataset of production-quality procedural materials and combine them with a constrained tree search inference algorithm that ensures syntactic validity while efficiently navigating the program space. Our experimental results show that our multimodal program synthesis method is more efficient in both unconditional and conditional graph synthesis with higher visual quality and fidelity than text-only baselines, establishing new state-of-the-art performance.
@@ 1 Introduction
Procedural materials have become increasingly important in modern 3D content creation, offering artists greater control and flexibility in designing surface appearances for digital assets. Unlike traditional image-based textures, which are constrained by fixed resolutions and limited editability, procedural material modeling tools like Adobe Substance Designer (Adobe, 2025c) or Blender (Blender, 2025) leverage node-based graphs to generate textures programmatically. This enables resolution-independent execution, high-level parametric control, and non-destructive editing workflows that have proven valuable in industries such as game development, film production, and VR/AR applications (Ebert et al., 2003). More specifically, a procedural material is defined as a directed graph where nodes represent texture generators (e.g., noise functions, patterns) or filtering operations (e.g., blurs, color adjustments), and edges encode the flow of data between these operations, ultimately producing the texture maps required by physically-based rendering (PBR) models (Pharr et al., 2016) (cf. Figure 1). However, the complexity of crafting these procedural material graphs presents a substantial barrier to entry, creating a pressing need for automated and semi-automated approaches to support material artists at all levels of proficiency.
@@ 2 Related Work
Our work builds upon recent advances in neural program synthesis (Parisotto et al., 2017; Devlin et al., 2017; Ellis et al., 2021). Traditional program synthesizers require formal specifications and employ search or logical derivation to produce programs that provably satisfy these specifications (Alur et al., 2013). Recently, large language models have demonstrated impressive capabilities in this domain (Huynh & Lin, 2025; Li et al., 2025b; Lozhkov et al., 2024; Li et al., 2023b; Rozière et al., 2023; Fried et al., 2023; Li et al., 2022; Chen et al., 2021). However, current research predominantly targets high-resource programming languages such as Python, Java, and JavaScript (Zan et al., 2023; Huynh & Lin, 2025). In contrast, our work synthesizes graphics programs, which pose unique challenges due to domain-specific requirements and considerable data scarcity, establishing it as a distinct research area.
@@
Deep learning approaches have shown strong performance in synthesizing graphics programs that compile to visual outputs (Ellis et al., 2018; 2019; Ganin et al., 2018). This progress has been accelerated by the emergence of large multimodal models, particularly vision-language models that bridge visual and textual domains (Alayrac et al., 2022; Liu et al., 2023; Belouadi et al., 2024b; Kulits et al., 2024; Li & Ellis, 2024; Kapur et al., 2025). The field encompasses both controlled experimental settings using domain-specific languages (Ellis et al., 2018; Tian et al., 2019; Sharma et al., 2018; Cámara et al., 2023; Kulits et al., 2024; Kapur et al., 2025) and practical applications. Notable examples include systems for generating scientific figures using TikZ (Belouadi et al., 2024a; b; 2025; Laurençon et al., 2024; 2024; Tong et al., 2024; Zhang et al., 2025) and automating data visualization (Mackinlay, 1986; Roth et al., 1994; Luo et al., 2021; Wu et al., 2024; Voigt et al., 2024). However, these approaches generate code designed for text-based editing and therefore do not face the unique circumstances of node graphs in procedural material synthesis that our work addresses.
@@ 6 Experiments
We build MultiMat models upon the QWen2.5VL (7B) foundation model (Bai et al., 2025). We train and evaluate separate models for unconditional generation (cf. §6.1) and inverse procedural material synthesis (cf. §6.2). Across all model variants, we maintain a consistent maximum sequence length of 8 192 tokens. The training setup consists of 5 epochs using a learning rate of and a batch size of 128. To ensure diversity in our generated outputs, we set the inference sampling parameters to a temperature of 0.8 and a top-p value of 0.95. We provide examples in Figure 5 and Appendix A.
@@ 6.1 Evaluation of Unconditional Generation
For text-only procedural material synthesis, VLMaterial represents the current state-of-the-art approach. However, its Blender-specific training makes direct comparison with our method difficult. We therefore create VLMaterial (SBS) by retraining a VLMaterial-style model on our dataset for fair comparison. Since VLMaterial (SBS) does not receive any images in the unconditional setting, we base it on the larger and more powerful text-only model QWen3 (8B; Yang et al., 2025a), giving it a slight advantage over our models. While graphics program synthesis research typically also benchmarks against proprietary large language models such as GPT-4o (OpenAI et al., 2024) or Claude 4 (Anthropic, 2025), which have demonstrated competitive performance in related domains (Belouadi et al., 2024a; b; 2025; Rodriguez et al., 2025), these models’ unfamiliarity with CompactSBS and inability to produce valid SBS output preclude their inclusion as baselines.
@@ 7 Analysis & Discussion
Our results show that model performance steadily improves as the degree of visualization of graphs increases, with MultiMat (Graph) achieving the best results overall. This finding aligns with how humans interact with procedural materials—through visual node graph interfaces—and validates established UX design principles in this domain. The qualitative examples in Figure 5 further illustrate this trend, with VLMaterial+ (SBS) struggling to generate faithful outputs, indicating that purely text-based approaches are not ideal for expressive node graph systems like Designer. This limitation persists even with more powerful base models, as our unconditional generation experiments confirm. Beyond architectural improvements, our tree search algorithm enables more efficient graph generation; without it, models may have to resort to sampling complete outputs for validation (the inference approach used by previous methods), which is expensive. For instance, disabling tree search causes NER of VLMaterial (SBS) to deteriorate further from 14.846 to 33.953, highlighting how our search strategy can improve inference without further training.
@@ 8 Conclusion
Although our models and baselines use the same or similar base models, they generate graphs in fundamentally different ways, resulting in considerable differences in training efficiency. Text-only models like VLMaterial can process entire graphs as single training examples, whereas MultiMat must adapt the visual context for each individual node, effectively processing training examples one node at a time. This difference leads to much longer training times: while VLMaterial completes training in a few hours on 80GB GPUs, MultiMat models require several days on the same hardware despite being trained on a comparable number of tokens. However, this training inefficiency does not affect inference, where both approaches achieve comparable generation speeds.
Acknowledgement
The authors wish to thank the team providing access to the Substance 3D Assets Repository, as well as the Substance Automation Python API. The work was done while the first author was an intern at Adobe Research in France.
@@ Appendix A Additional Examples
Figure 6 complements Figure 5 by showcasing failure cases where our models struggle to produce faithful outputs, though notably, the outputs from MultiMat+ (Graph) and MultiMat+ (Mixed) still demonstrate superior representation of the input compared to VLMaterial (SBS). Beyond these conditional generation examples, Figure 7 presents unconditional samples generated by MultiMat (Graph), which exhibit high visual quality with realistic material properties. Adjacent to these rendered materials, we visualize their underlying material graphs in the same format used as model input. In Figure 8, we show a graph in CompactSBS representation to give an impression of the structure of our format.

### v2

# testphase=new-or-1 MultiMat: Multimodal Program Synthesis for Procedural Materials using Large Multimodal Models

@@ Abstract
Material node graphs are programs that generate the 2D channels of procedural materials, including geometry such as roughness and displacement maps, and reflectance such as albedo and conductivity maps. They are essential in computer graphics for representing the appearance of virtual 3D objects parametrically and at arbitrary resolution. In particular, their directed acyclic graph structure and intermediate states enable a modular, interpretable workflow for interactive appearance modeling. However, creating such graphs remains challenging and typically requires professional training. While recent neural program synthesis approaches attempt to simplify this process, they solely represent graphs as textual programs, failing to capture the inherently visual-spatial nature of node graphs that makes them accessible to humans. To address this gap, we present MultiMat, a multimodal program synthesis framework that leverages large multimodal models to process both visual and textual graph representations for improved generation of procedural material graphs. We train our models on a new dataset of production-quality procedural materials and combine them with a constrained tree search inference algorithm that ensures static correctness while efficiently navigating the program space. Our experimental results show that our multimodal program synthesis method is more efficient in both unconditional and conditional graph synthesis with higher visual quality and fidelity than text-only baselines, establishing new state-of-the-art performance.
@@ 1 Introduction
Procedural materials have become increasingly important in modern 3D content creation, offering artists greater control and flexibility in designing surface appearances for digital assets. Unlike traditional image-based textures, which are constrained by fixed resolutions and limited editability, procedural material modeling tools like Adobe Substance Designer (Adobe, 2025c) or Blender (Blender, 2025) leverage node-based graphs to generate textures programmatically. This enables resolution-independent execution, high-level parametric control, and non-destructive editing workflows that have proven valuable in industries such as game development, film production, and VR/AR applications (Musgrave et al., 2002). More specifically, a procedural material is defined as a directed graph where nodes represent texture generators (e.g., noise functions, patterns) or filtering operations (e.g., blurs, color adjustments), and edges encode the flow of data between these operations, ultimately producing the texture maps required by physically-based rendering (PBR) models (Pharr et al., 2016) (cf. Figure 1). However, the complexity of crafting these procedural material graphs presents a substantial barrier to entry, creating a pressing need for automated and semi-automated approaches to support material artists at all levels of proficiency.
@@ 2 Related Work
Our work builds upon recent advances in neural program synthesis (Parisotto et al., 2017; Devlin et al., 2017; Thakoor et al., 2018; Ye et al., 2021; Ellis et al., 2021). Traditional program synthesizers require formal specifications and employ search or logical derivation to produce programs that provably satisfy these specifications (Alur et al., 2013). Recently, large language models have demonstrated impressive capabilities in this domain (Huynh & Lin, 2025; Li et al., 2025b; Lozhkov et al., 2024; Li et al., 2023b; Rozière et al., 2023; Fried et al., 2023; Li et al., 2022; Chen et al., 2021). However, current research predominantly targets high-resource programming languages such as Python, Java, or JavaScript (Zan et al., 2023; Huynh & Lin, 2025). In contrast, our work synthesizes graphics programs, which pose unique challenges due to domain-specific requirements and considerable data scarcity, establishing it as a distinct research area.
@@
Deep learning approaches have shown strong performance in synthesizing graphics programs that compile to visual outputs (Ellis et al., 2018; 2019; Ganin et al., 2018). This progress has been accelerated by the emergence of large multimodal models, particularly vision-language models that bridge visual and textual domains (Alayrac et al., 2022; Liu et al., 2023; Belouadi et al., 2024b; Kulits et al., 2024; Li & Ellis, 2024; Kapur et al., 2025; Lin et al., 2025; Xu et al., 2025). The field encompasses both controlled experimental settings using domain-specific languages (Ellis et al., 2018; Tian et al., 2019; Sharma et al., 2018; Cámara et al., 2023; Kulits et al., 2024; Kapur et al., 2025; Wen et al., 2025) and practical applications. Notable examples include systems for generating scientific figures using TikZ (Belouadi et al., 2024a; b; 2025; Laurençon et al., 2024; 2024; Tong et al., 2024; Zhang et al., 2025) and automating data visualization (Mackinlay, 1986; Roth et al., 1994; Luo et al., 2021; Wu et al., 2024; Voigt et al., 2024). However, these approaches generate code designed for text-based editing and therefore do not face the unique circumstances of node graphs in procedural material synthesis that our work addresses.
@@ 4.1 Multimodal Program Synthesis
At their core, both approaches remain autoregressive language models, and MultiMat can be trained by minimizing a cross-entropy objective:
where is the expected token of in our intermediate text format at position , represents previous tokens, denotes the input conditions (which can be empty for unconditional generation), and represents the model parameters.
@@ 6 Experiments
We build MultiMat models upon the QWen2.5VL (7B) vision-language model which leverages a late fusion approach to combine image and text tokens (Bai et al., 2025). We train and evaluate separate models for unconditional generation (cf. §6.1) and inverse procedural material synthesis (cf. §6.2). We also conduct a human evaluation (cf. §6.3). Across all model variants, we maintain a consistent maximum sequence length of 8 192 tokens. The training setup consists of 5 epochs using AdamW (Loshchilov & Hutter, 2019), a learning rate of , and a batch size of 128. To ensure diversity in our generated outputs, we set the inference sampling parameters to a temperature of 0.8 and a top-p value of 0.95. We provide examples in Figure 6 and Appendix A. We ablate incremental tree search in §7.
@@ 6.1 Evaluation of Unconditional Generation
For text-only procedural material synthesis, VLMaterial represents the current state-of-the-art approach. However, its Blender-specific training makes direct comparison with our method difficult. We therefore create VLMaterial (SBS) by retraining a VLMaterial-style model on our dataset for fair comparison. Unlike the objective in equation (1), VLMaterial is trained to generate complete graphs in a single pass. However, during inference, we can still validate nodes as they are generated and roll back upon detecting irreparable errors (cf. §4.2) or repair them after generation completes (cf. §4.3). This means the progression from VLMaterial (SBS) to MultiMat (Mixed) to MultiMat (Graph) represents a comparable, gradual shift from complete graph generation toward iterative node generation. Since VLMaterial (SBS) does not receive any images in the unconditional setting, we base it on the larger and more powerful text-only model QWen3 (8B; Yang et al., 2025a), giving it a slight advantage over our models. While graphics program synthesis research typically also benchmarks against proprietary large language models such as GPT-4o (OpenAI et al., 2024) or Claude 4 (Anthropic, 2025), which have demonstrated competitive performance in related domains (Belouadi et al., 2024a; b; 2025; Rodriguez et al., 2025), these models’ unfamiliarity with CompactSBS and inability to produce valid SBS output preclude their inclusion as baselines.
@@ Results
6.3 Human Evaluation
To corroborate our automatic evaluation results, we conduct a human evaluation. We employ comparative annotation (Thurstone, 1927) and focus on the image reconstruction/inverse rendering use case, which allows for intuitive human assessment (cf. Figure 6). Annotators receive triplets of rendered generated materials from VLMaterial+ (SBS), MultiMat (Mixed), and MultiMat (Graph) and identify which output best and least resembles the input image. Following Hu et al. (2023); Li et al. (2025a), we generate multiple programs () per model and image, selecting the result with the highest DreamSim score as the final candidate. We test 33 input materials from graphs filtered during preprocessing (e.g., due to excessive length), which represent particularly challenging cases. Eight expert annotators with extensive procedural material experience assess each triplet in randomized order. As shown in Figure 5, annotators rank VLMaterial+ (SBS) considerably lower than MultiMat+ (Mixed) and MultiMat+ (Graph), and prefer MultiMat+ (Graph) over MultiMat+ (Mixed). These findings align with our automatic evaluation rankings and demonstrate our approach’s effectiveness in generating perceptually similar materials.
@@ 7 Analysis & Discussion
Our comparisons demonstrate that model performance improves steadily as the degree of graph visualization increases, with MultiMat (Graph) achieving the highest performance overall (cf. Tables 2 & 3; Figure 5). This finding aligns with how humans interact with procedural materials—through visual node graph interfaces—and validates established UX design principles in this domain.
| Models | Deletion↓ | Conversion↓ |
| VLMaterial (SBS) | ||
| MultiMat (Mixed) | ||
| MultiMat (Graph) |
The qualitative examples in Figure 6 further illustrate this trend, with VLMaterial+ (SBS) struggling to generate faithful outputs, indicating that purely text-based approaches are not ideal for expressive node graph systems like Designer. This limitation persists even with more powerful base models, as our unconditional generation experiments confirm. Beyond architectural improvements, our tree search algorithm enables more efficient graph generation; without it, models may have to resort to sampling complete outputs for validation (the inference approach used by previous methods), which is expensive. For instance, disabling tree search causes NER of VLMaterial (SBS) to deteriorate further from 14.846 to 33.953, highlighting how our search strategy can improve inference without further training. The impact of automatic error repair is more nuanced, as shown in Table 4. Only approximately 1% of nodes generated by MultiMat contain hallucinated parameters, and fewer than 6.5% require conversion. In contrast, VLMaterial exhibits nearly double the scores for both repair mechanisms. This difference demonstrates that VLMaterial requires considerably more repair than our models and supports our claim that our models possess a better understanding of graph structures. Notably, since corrections are not fed back to the models, these results reflect their intrinsic generation capabilities.
@@ 8 Conclusion
Although our models and baselines use the same or similar base models, they generate graphs in fundamentally different ways, resulting in considerable differences in training efficiency. Text-only models like VLMaterial can process entire graphs as single training examples, whereas MultiMat must adapt the visual context for each individual node, effectively processing training examples one node at a time. This difference leads to much longer training times: while VLMaterial completes training in a few hours on 80GB GPUs, MultiMat models require several days on the same hardware despite being trained on a comparable number of tokens. Nevertheless, since the amount of procedural materials is very small (regardless of the dataset), training times remain within acceptable bounds in absolute terms, despite the relative differences between methods. Additionally, both approaches achieve a more similar throughput during inference.
@@
Acknowledgements
We thank the Adobe Substance 3D team for providing access to the Substance 3D Assets Repository and the Substance Automation Python API. We also thank our annotators for their valuable time. This work was conducted while the first author was an intern at Adobe Research, France.
@@ Appendix A Additional Examples
In Figure 7 we provide additional qualitative examples. MultiMat+ (Mixed) consistently surpasses VLMaterial+ (SBS), while MultiMat+ (Graph) demonstrates the strongest results overall. Figure 8 complements Figures 6 & 7 by showcasing failure cases where our models struggle to produce faithful outputs, though notably, the outputs from MultiMat+ (Graph) and MultiMat+ (Mixed) still demonstrate superior representation of the input compared to VLMaterial (SBS). Beyond these conditional generation examples, Figure 9 presents unconditional samples generated by MultiMat (Graph), which exhibit high visual quality with realistic material properties. Adjacent to these rendered materials, we visualize their underlying material graphs in the same format used as model input. In Figure 10, we show a graph in CompactSBS representation to give an impression of the structure of our format.
| Input | VLMaterial+ (SBS) | MultiMat+ (Mixed) | MultiMat+ (Graph) |